# A functional shunt in the umbilical cord: the role of coiling in solute and heat transfer

### v1

@@ III.2 Configurations yielding maximum and minimum exchange
Figure 3: Solute exchange flux $N$ for different cord configurations in the ($\theta_{1}$, $\Delta\theta$) parameter space (see Fig. 1b). The geometrical parameters used were $r_{v}=0.25,r_{a}=0.12,d=0.04,\eta=0,\Omega=3.77$ for (a) $R_{v}=0$ (a straight vein, as shown in (c)) and (b) $R_{v}=0.21$. Cord configurations predicted by the computational model (d) ($\theta_{1}=0.9\pi$; $\Delta\theta=0.2\pi$) that minimise and (e) ($\theta_{1}=0.5\pi$; $\Delta\theta=\pi$) maximise solute exchange between the vessels (shown in (b)). Solute concentrations are plotted in normalised units, varying from zero on the arterial surfaces (blue) to one on the surface of the vein (red). (f) Magnification of a region in (b) showing estimated angles obtained from histology and from ultrasound images of human cord cross-sections from healthy pregnancies. The line of symmetry (dashed) has a slope of $-2$ and a zero intercept at $\theta_{1}=\pi$. (g) and (h) are examples of histology images with irregular shapes, circularity $<1$ ([25], reproduced under CC BY 4.0, and [21], reproduced by permission of Taylor & Francis Ltd, tandfonline.com). (i) Ultrasound image of a cord with approximately unit circularity ([24], reproduced by permission of Karger Publishers, © 2009).
@@
We now consider cross-sectional cord configurations with two arteries in different positions by changing the angles $\theta_{1}$ and $\Delta\theta=|\theta_{2}-\theta_{1}|$ (Fig. 1b). We keep the artery-vein distance ($d=0.04$) constant and use $\theta_{1}\in[0,\pi]$, while $\theta_{2}\in[0,2\pi)$. We plot the exchange flux $N$ for all combinations of $\theta_{1}$ and $\theta_{2}$ for tightly coiled ($\Omega=3.77$) configurations with a straight vein and zero offset $R_{v}=0$ (Fig. 3a) and for a coiled vein with an offset from the cord centreline $R_{v}=0.25$ (Fig. 3b). Configurations with overlapping arteries were excluded from the map; the smallest possible difference in angles such that the arteries do not intersect was $\Delta\theta=0.2\pi$.
@@
When the vein is located at the centre of the cord ($R_{v}=0$; Fig. 3c), the exchange flux is invariant with respect to $\theta_{1}$ and depends only on the relative position $\Delta\theta$ of the two UAs around the UV (Fig. 3a). The exchange is maximised at $\Delta\theta=\pi$ when the arteries are aligned on opposite sides of the vein (Fig. 3c). Increasing the offset of the vein from the cord’s centre ($R_{v}>0$), and hence twisting the vein, amplifies the solute flux (the maximum $N$ increases from about 15 to 26; see Fig. 3b). Furthermore, the displaced vein, which acts as a source, breaks the radial symmetry of the concentration field and produces a complex landscape of enhanced and reduced values for the exchange flux, as shown in Fig. 3(b). The region around the global maximum ($\theta_{1}\approx 0.5\pi$ and $\Delta\theta\approx\pi$) corresponds to the configuration where the arteries are aligned on either side of the vein, perpendicular to the $x$-axis that passes through the centre of the cord and centre of the vein (Fig. 3d). Similarly, local maxima are attained when the arteries are close to each other in the region $\theta_{1}\in[\pi/4,\pi/2]$. The global minimum ($\theta_{1}\approx 0.9\pi$ and $\Delta\theta\approx 0.2\pi$) is achieved when arteries are close to each other and are placed along the $x$-axis near the centre of the cord (Fig. 3e).
@@
The configurations that maximise the exchange flux in both scenarios (Fig. 3c,e) are associated with UAs that coil around the UV, allowing for a larger surface area. However, a more general map in Fig. 3(b) reveals that the exchange flux is smaller when the two UAs are aligned close to the $x$-axis ($\theta_{1}\approx\pi$) than when they are aligned orthogonally to it ($\theta_{1}\approx\pi/2$). This indicates the importance of a specific three-dimensional arrangement of the cord’s vessels, which goes beyond the effective surface area of the UA-UV interface.
@@
We hypothesise that the enhanced exchange between helical vessels might boost the supply of oxygen to the walls of the umbilical arteries and hence function as a ‘virtual’ vasa vasorum. In Fig. 4(a) we show the effects of helicity $\Omega$, and the relative solute concentration difference between the UA and the UV, $\beta$ on the extravascular uptake flux $N_{u}$ (see (2)). Similarly to Fig. 2, the uptake flux increases with coiling; however, for weak tissue metabolism, the effect is less noticeable. Fig. 4(b) and (c) show similar effects to Fig. 2(a), where concentration fields become more uniform as $\Omega$ increases. For higher metabolic rate $\alpha$, the impact of helicity gets stronger, and the distortions in the concentration fields in the cord tissue also have a stronger impact on the uptake flux. Fig. 4(d) shows boundary layers around the arteries, which have an elliptical shape due to the effects of helicity on the concentration field. Additionally, the value of $\alpha$ for which the maximum flux is attained at increases for larger $\Omega$.
@@
Fig. 4(a) shows a non-monotonic increase in the extravascular uptake flux as the metabolic rate parameter increases. Clearly, for $\alpha=0$, the uptake flux in the cord tissue must be zero. Large $\alpha$ leads to significant spatial gradients of concentration localised in narrow regions around the vessels (Fig. 4d), reducing solute exchange between the UA and the UV. We assess the impact of metabolism on the oxygen supply to the avascular cord tissue by excluding thin strips of size $h$ (comparable to the vessel wall thickness). Therefore, the uptake flux $N_{u}$ (Fig. 4a) tends to zero for sufficiently strong metabolism ($\alpha\gg 1/h^{2}$), when the tissue oxygenation outside the immediate vicinity of the vascular lumen is negligible. Fig. 4(a) also illustrates the role of the relative venous–arterial concentration difference $\beta$ for the baseline case of $\beta=1$ (oxygen concentration in the UV is twice that of the UA) and for $\beta=2$ (the UV concentration is three times that of the UA). The apparent drop in peak $N_{u}$ at higher $\beta$ masks an increase in the absolute uptake flux that scales with this concentration difference $C_{v}-C_{a}$ (in dimensional variables; see (2)).
@@
To explore the contribution of the umbilical vein to cord tissue oxygenation more directly, we consider a base case in which oxygen in the cord tissue is supplied only by the UA. We compare the net uptake flux $N_{u}$ to this base uptake flux $N_{b}$ by calculating $N_{s}=(N_{u}-N_{b})/N_{b}$, which quantifies the additional oxygen supplied to the tissue due to the presence of the UV (see the Supplement S4 for more details). Fig. 4(e) shows how $N_{s}$ varies with $\alpha$ for different vessel separation $d$. Similarly to Fig. 4(a) there is a non-monotonic increase in excess oxygen supply. As vessels get closer to each other, a large $\alpha$ similarly blocks the interactions between the umbilical vessels and hence diminishes the exchange between them. However, when vessel separation $d$ is small the additional oxygen supply to the cord tissue is twice as large compared to the base case.
@@ IV Discussion
In this study, we examine the hypothesis that there exists a functional shunt between the umbilical vessels through a mathematical model that describes solute diffusion and uptake in the cord tissue. By exploiting a natural transformation of the coordinate system, we encapsulate the complex three-dimensional helical geometry of the umbilical vessels in a computationally efficient and interpretable two-dimensional problem, which is formulated in the cross-sectional plane of the cord. The model reveals that cord coiling and vascular proximity promote solute exchange between the umbilical vessels (Fig. 2). In the limit of moderate helicity ($\Omega\leq 1$), the model shows that the exchange flux is approximately proportional to the square of the cord helicity, or equivalently the UCI augmented by the cord diameter $D$:  $\Omega^{2}\sim{\color[rgb]{0,0,0}D}^{2}(\text{UCI})^{2}$. Similarly, in the limit of a thin gap between the umbilical vessels, separated by the relative distance $d\ll 1$, the model predicts that the dominant contribution to the exchange flux is $\mathcal{O}(d^{-1/2})$. Similar to the earlier results for blood flow modelling in the umbilical cord [13], these geometric determinants show that the UCI alone does not capture the essential features of solute exchange and thus can be a poor indicator of clinical pathology.
@@
The simulations also reveal a non-trivial relationship between the vascular configurations in the cross-section of the umbilical cord and the extent of solute exchange between the vessels (see Fig. 3). We use theoretical predictions to interpret the structural images of cords from healthy pregnancies. In all the samples we investigated the umbilical arteries were positioned approximately symmetrically with respect to the axis, passing through the cord centre and the centre of the vein, and that the observed configurations tend to minimise shunting between the umbilical vessels. To put this in the context of the typical transported solutes, we assess the flux exchanged between the umbilical vessels relative to the flux delivered to the fetus by the umbilical vein, as quantified by a Damköhler number $\mathrm{Da}$ (3) (see Supplement S8). Oxygen transport is diffusion-limited (advection-dominated) in the umbilical cord, (as represented by a small $\mathrm{Da}$), implying that the amount of oxygen exchanged between the cord vessels can be neglected. In contrast, the diffusive coupling for heat transfer may be significant, because the relevant $\mathrm{Da}$ is significantly larger. Given the tendency to minimise this coupling in the measured cord configurations, it opens an intriguing possibility that the cord structure can be partially explained by the demands of fetal thermoregulation.
@@
Although the human umbilical cord tissue lacks vasa vasorum [27], the cord’s WJ contains mesenchymal stem cells (MSCs), fibroblasts, and other connective tissue cells [18], all of which are metabolically active and depend on oxygen to perform their functions, which are essential for the health of the cord (such as tissue repair by the MSCs and collagen synthesis by the fibroblasts). Our model demonstrates how the functional shunt between the umbilical vessels can act as ‘virtual’ vasa vasorum by amplifying the oxygen supply to the cord tissue. While the cells residing in the WJ are likely adapted to a relatively low oxygen environment compared to other tissues [27], we hypothesise that abnormal coiling and vascular proximity could adversely affect the development and function of the cord tissue.
@@
Our modelling approach has many limitations and assumptions that warrant further study. For example, we neglected the effects of potential solute heterogeneity on the surface of umbilical vessels (which is important in parallel uncoiled exchanger systems [28]). However, the length of the cord and the helicity-induced secondary flows [29, 30] will promote the mixing of solutes in umbilical vessels, justifying to some extent the azimuthally uniform approximation. Future work can generalise this model by considering non-zero axial concentration gradients to study exchange along the whole cord and explore more accurately how the functional shunt affects the transport of solutes between the placenta and the fetus.
@@
Additionally, in this study, we primarily analysed the cord structure based on histological images; however, the in vivo structure of the cord is prone to significant perturbations post-delivery. For example, the umbilical vein lumen collapses due to changes in pressure, while the umbilical artery lumen contracts and clots within in help avoid neonatal blood loss [31]. These effects can be seen in Fig. 3(g,h), where the vein has an elliptical shape, the arterial lumens are not readily visible, and the cord’s cross-section can have an irregular (non-circular) shape. In contrast, the ultrasound images better reproduce the in utero environment since the vessels are not collapsed or constricted, with more circular shapes and visible lumens (Fig. 3i). Our analysis relies on the relative centre-of-mass locations of the umbilical vessels, rather than finer shape details, making our results less sensitive to these ex vivo distortions. Nevertheless, in future studies, ultrasound images should be used for a more accurate representation of the umbilical vascular structure.
@@
We have also limited our analysis to the cords from normal human pregnancies. Pathological studies often report thin cords with a significant decrease in WJ [32], and instances of cords with a single UA, which have been associated with fetal growth restriction, pre-eclampsia and other adverse pregnancy outcomes  [33]. Similarly, the morphology of the umbilical cord is known to vary significantly across different placental mammals [19]. These structural differences motivate the need for a systematic quantitative comparison of the cord’s performance in healthy and pathological human pregnancies, as well as its interspecies variations [34, 35].
@@
Data Accessibility. All data needed to evaluate the conclusions in the paper are present in the paper and/or the Supplementary Materials. Additional data related to this paper may be requested from the authors.
@@
Funding. This work was supported by the A*STAR Scholarship Award to TW. EDJ, OEJ & ILC also gratefully acknowledge partial support by the Wellcome Leap In Utero Programme.

### v3

@@ III.2 Configurations yielding maximum and minimum exchange
Figure 3: Solute exchange flux $N$ for different cord configurations in the ($\theta_{1}$, $\Delta\theta$) parameter space (see Fig. 1b). The geometrical parameters used were $r_{v}=0.25,r_{a}=0.12,d=0.04,\eta=0,\Omega=3.77$ (see Table 1) for (a) $R_{v}=0$ (a straight vein, as shown in (c)) and (b) $R_{v}=0.21$. Cord configurations predicted by the computational model (e) ($\theta_{1}=0.9\pi$; $\Delta\theta=0.2\pi$) that minimise and (d) ($\theta_{1}=0.5\pi$; $\Delta\theta=\pi$) maximise solute exchange between the vessels (shown in (b)). Solute concentrations are plotted in normalised units, varying from $C=0$ on the arterial surfaces (blue) to $C=1$ on the surface of the vein (red). (f) Magnification of a region in (b) showing estimated angles obtained from histology and from ultrasound images of human cord cross-sections from healthy pregnancies. The line of symmetry (dashed) has a slope of 2-2 and a zero intercept at $\theta_{1}=\pi$. (g) and (h) are examples of histology images with irregular shapes, circularity <1<1 ([30], reproduced under $C$ BY 4.0, and [26], reproduced by permission of Taylor & Francis Ltd, tandfonline.com). (i) Ultrasound image of a cord with approximately unit circularity ([29], reproduced by permission of Karger Publishers, © 2009).
@@
We now consider cross-sectional cord configurations with two arteries in different positions by changing the angles $\theta_{1}$ and $\Delta\theta=|\theta_{2}-\theta_{1}|$ (Fig. 1b). We keep the artery-vein distance ($d=0.04$) constant and use $\theta_{1}\in[0,\pi]$, while $\theta_{2}\in[0,2\pi)$. We plot the exchange flux $N$ for all combinations of $\theta_{1}$ and $\theta_{2}$ for tightly coiled ($\Omega=3.77$) configurations with a straight vein and zero offset $R_{v}=0$ (Fig. 3a,c) and for a coiled vein with an offset from the cord centreline $R_{v}=0.25$ (Fig. 3b,d). Configurations with overlapping arteries were excluded from the map; the smallest possible difference in angles such that the arteries do not intersect was $\Delta\theta=0.2\pi$.
@@
When the vein is located at the centre of the cord ($R_{v}=0$; Fig. 3c), the exchange flux is invariant with respect to $\theta_{1}$ and depends only on the relative position $\Delta\theta$ of the two UAs around the UV (Fig. 3a). The exchange is maximised at $\Delta\theta=\pi$ when the arteries are aligned on opposite sides of the vein (Fig. 3c). Increasing the offset (helical radius) of the vein from the cord’s centre ($R_{v}>0$), and hence twisting the vein, amplifies the solute flux (the maximum $N$ increases from about 15 to 26; see Fig. 3b). Furthermore, the displaced vein, which acts as a source, breaks the radial symmetry of the concentration field and produces a complex landscape of enhanced and reduced values for the exchange flux, as shown in Fig. 3(b). The region around the global maximum ($\theta_{1}\approx 0.5\pi$ and $\Delta\theta\approx\pi$) corresponds to the configuration where the arteries are aligned on either side of the vein, perpendicular to the $x$-axis that passes through the centre of the cord and centre of the vein (Fig. 3d). Similarly, local maxima are attained when the arteries are close to each other in the region $\theta_{1}\in[\pi/4,\pi/2]$. The global minimum ($\theta_{1}\approx 0.9\pi$ and $\Delta\theta\approx 0.2\pi$) is achieved when arteries are close to each other and are placed along the $x$-axis near the centre of the cord (Fig. 3e).
@@
The configurations that maximise the exchange flux in both scenarios (Fig. 3c,d) are associated with UAs that straddle the UV, allowing for a larger surface area. However, a more general map in Fig. 3(b) reveals that the exchange flux is smaller when the two UAs are aligned close to the $x$-axis ($\theta_{1}\approx\pi$;  Fig. 3e) than when they are aligned orthogonally to it ($\theta_{1}\approx\pi/2$). This indicates the importance of a specific three-dimensional arrangement of the cord’s vessels, which goes beyond the effective surface area of the UA-UV interface.
@@
We hypothesise that the enhanced exchange between helical vessels might boost the supply of oxygen to the walls of the umbilical arteries and hence function as a ‘virtual’ vasa vasorum. In Fig. 4(a) we show the effects of helicity $\Omega$, and the relative solute concentration difference between the UA and the UV $\beta$, on the extravascular uptake flux $N_{u}$ (see (2)). Similarly to Fig. 2, the uptake flux increases with coiling; however, for weak tissue metabolism, the effect is less noticeable. Fig. 4(b) and (c) show similar effects to Fig. 2(a), where concentration fields become more uniform as $\Omega$ increases. For higher metabolic rate $\alpha$, the impact of helicity gets stronger, and the distortions in the concentration fields in the cord tissue also have a stronger impact on the uptake flux. Fig. 4(d) shows boundary layers around the arteries, which have an elliptical shape due to the effects of helicity on the concentration field. Additionally, the value of $\alpha$ for which the maximum flux is attained increases for larger $\Omega$.
@@
Fig. 4(a) shows a non-monotonic increase in the extravascular uptake flux $N_{u}$ (2) as the metabolic rate parameter increases. Clearly, for $\alpha=0$, the uptake flux in the cord tissue must be zero. Large $\alpha$ leads to significant spatial gradients of concentration localised in narrow regions around the vessels (Fig. 4d), reducing solute exchange between the UA and the UV. We assess the impact of metabolism on the oxygen supply to the avascular cord tissue by excluding thin strips of size $h$ (comparable to the vessel wall thickness). Therefore, the uptake flux $N_{u}$ (Fig. 4a) tends to zero for sufficiently strong metabolism ($\alpha\gg 1/h^{2}$), when the tissue oxygenation outside the immediate vicinity of the vascular lumen is negligible. Fig. 4(a) also illustrates the role of the relative venous–arterial concentration difference $\beta$ for the baseline case of $\beta=1$ (oxygen concentration in the UV is twice that of the UA) and for $\beta=2$ (the UV concentration is three times that of the UA). The apparent drop in peak $N_{u}$ at higher $\beta$ masks an increase in the absolute uptake flux that scales with this concentration difference $C_{v}-C_{a}$ (in dimensional variables; see (2)).
@@
To explore the contribution of the umbilical vein to cord tissue oxygenation more directly, we consider a base case in which oxygen in the cord tissue is supplied only by the UA. We compare the net uptake flux $N_{u}$ to this base uptake flux $N_{b}$ by calculating $N_{s}=(N_{u}-N_{b})/N_{b}$, which quantifies the additional oxygen supplied to the tissue due to the presence of the UV (see the Supplement S4 for more details). Fig. 4(e) shows how $N_{s}$ varies with $\alpha$ for different vessel separations $d$. Similarly to Fig. 4(a) there is a non-monotonic increase in excess oxygen supply. As vessels get closer to each other, a large $\alpha$ similarly blocks the interactions between the umbilical vessels and hence diminishes the exchange between them. However, when the vessel separation $d$ is small the additional oxygen supply to the cord tissue is almost twice as large compared to the base case (see Fig. 4e).
@@ IV Discussion
In this study, we examine the hypothesis that there exists a functional shunt between the umbilical vessels through a mathematical model that describes solute diffusion and uptake in the cord tissue. By exploiting a natural transformation of the coordinate system, we encapsulate the complex three-dimensional helical geometry of the umbilical vessels in a computationally efficient and interpretable two-dimensional problem, which is formulated in the cross-sectional plane of the cord. The model reveals that cord coiling and vascular proximity promote solute exchange between the umbilical vessels (Fig. 2). In the limit of moderate helicity ($\Omega\leq 1$), the model shows that the exchange flux is approximately proportional to the square of the cord helicity, or equivalently the UCI augmented by the cord diameter $D$:  $\Omega^{2}\sim D^{2}(\text{UCI})^{2}$. Similarly, in the limit of a thin gap between the umbilical vessels, separated by the relative distance $d\ll 1$, the model predicts that the dominant contribution to the exchange flux is $\mathcal{O}(d^{-1/2})$. Similar to the earlier results for blood flow modelling in the umbilical cord [13, 32], these geometric determinants show that the UCI alone does not capture the essential features of solute exchange and thus can be a poor indicator of clinical pathology.
@@
The simulations also reveal a non-trivial relationship between the vascular configurations in the cross-section of the umbilical cord and the extent of solute exchange between the vessels (see Fig. 3). We use theoretical predictions to interpret the structural images of cords from healthy pregnancies. In all the samples we investigated, the umbilical arteries were positioned approximately symmetrically with respect to the axis, passing through the cord centre and the centre of the vein, so that the observed configurations tend to minimise shunting between the umbilical vessels. To put this in the context of the typical transported solutes, we assessed the flux exchanged between the umbilical vessels relative to the flux delivered to the fetus by the umbilical vein, as quantified by a Damköhler number $\mathrm{Da}$ (see (3)). Oxygen transport in the umbilical cord is diffusion-limited (advection-dominated, as represented by a small $\mathrm{Da}$), implying that the amount of oxygen exchanged between the cord vessels can be neglected. In contrast, the diffusive coupling for heat transfer may be significant, because the relevant $\mathrm{Da}$ is significantly larger. Given the tendency to minimise this coupling in the measured cord configurations, it opens an intriguing possibility that the cord structure can be partially explained by the demands of fetal thermoregulation.
@@
Although the human umbilical cord tissue lacks vasa vasorum [33], the cord’s WJ contains mesenchymal stem cells (MSCs), fibroblasts, and other connective tissue cells [18], all of which are metabolically active and depend on oxygen to perform their functions, which are essential for the health of the cord (such as tissue repair by the MSCs and collagen synthesis by the fibroblasts). Our model demonstrates how the functional shunt between the umbilical vessels can act as ‘virtual’ vasa vasorum by amplifying the oxygen supply to the cord tissue. While the cells residing in the WJ are likely adapted to a relatively low oxygen environment compared to other tissues [33], we hypothesise that abnormal coiling and vascular separation distance could adversely affect the development and function of the perivascular cord tissue; however, further experimental studies are needed to evaluate the physiological significance of this mechanism.
@@
Our modelling approach has many limitations and assumptions that warrant further study. For example, we neglected the effects of potential solute heterogeneity on the surface of umbilical vessels (which is important in parallel uncoiled exchanger systems [34]). However, the length of the cord and the helicity-induced secondary flows [35, 36] will promote the mixing of solutes in umbilical vessels, justifying to some extent the azimuthally uniform approximation. We also neglected axial concentration gradients and assumed a constant coiling pitch, whereas human cords can exhibit slow pitch variations [13]. Future work can generalise this model by considering non-zero axial concentration gradients to study exchange along the whole cord and explore more accurately how the functional shunt affects the transport of solutes between the placenta and the fetus.
@@
Additionally, in this study, we primarily analysed the cord structure based on histological images; however, the in vivo structure of the cord is prone to significant perturbations post-delivery. For example, the umbilical vein lumen collapses due to changes in pressure, while the umbilical artery lumen contracts and clots within it help avoid neonatal blood loss [37]. These effects can be seen in Fig. 3(g,h), where the vein has an elliptical shape, the arterial lumens are not readily visible, and the cord’s cross-section can have an irregular (non-circular) shape. In contrast, ultrasound images better reproduce the in utero environment since the vessels are not collapsed or constricted, with more circular shapes and visible lumens (Fig. 3i). Our analysis relies on the relative centre-of-mass locations of the umbilical vessels, rather than finer shape details, making our results less sensitive to these ex vivo distortions. Nevertheless, in future studies, ultrasound images should be used for a more accurate representation of the umbilical vascular structure.
@@
We have also limited our analysis to cords from normal human pregnancies. Pathological studies often report thin cords with a significant decrease in WJ [38], and instances of cords with a single UA, which have been associated with fetal growth restriction, pre-eclampsia and other adverse pregnancy outcomes [39]. Similarly, the morphology of the umbilical cord is known to vary significantly across different placental mammals [19]. These structural differences motivate the need for a systematic quantitative comparison of the cord’s performance in healthy and pathological human pregnancies (using multiple imaging modalities [40]), as well as its interspecies variations [41, 42].
@@
Data Accessibility. All data needed to evaluate the conclusions in the paper are present in the paper and/or the Supplementary Materials. The associated structural datasets and computational codes can be accessed via the Figshare repository https://doi.org/10.6084/m9.figshare.28382012.
@@
Funding. This work was supported by an A*STAR Scholarship Award to TW. EDJ, OEJ & ILC also gratefully acknowledge partial support by the Wellcome Leap In Utero Programme.